OpenStudy (anonymous):

Amy drove her car 253 miles. The car used 10.07 gallons of gasoline on the drive. Estimate the car's gasoline mileage in miles per gallon

OpenStudy (anonymous):

about 25 miles per gallon estimated

OpenStudy (tkhunny):

The answer is in the question. It asks for "miles per gallon". \(\dfrac{miles}{gallon}\longrightarrow\dfrac{253\;miles}{10.07\;gallon} = 25.12\;mpg\) This is he exact answer. It's really up to you how to "estimate" this. You MIGHT say, 253 isn't too far from 250. You MIGHT say, 10.07 isn't too far from 10. 250/10 = 25 and this reproduces the estimate given with no work shown.